Translate

New Live Courses

UpComing SoftWare Training Demos

๐Ÿš€ *Playwright With AI Testing Demo In Telugu*
๐Ÿ—“️ *Date:* 28 Oct2026, 08:00 PM IST
๐Ÿ“ *Register Now!:*
https://www.vlrt.in/3r
๐Ÿ‘ฅ *Join WA Community:*
https://www.vlrt.in/3w
๐Ÿ’ก *Course Content:*
https://www.vlrt.in/3c
▶️ *Demo Videos:*
https://www.vlrt.in/3v

๐Ÿš€ *Gen AI Engineer (Production Focused)*
๐Ÿ—“️ *Date:* 29th Oct 2026, 06:00PM IST
๐Ÿ“ *Register Now!:*
https://www.vlrt.in/gr
๐Ÿ‘ฅ *Join WA Community:*
https://www.vlrt.in/gw
๐Ÿ’ก*Course Content:*
https://www.vlrt.in/ai
▶️ *Demo Videos:*
https://www.vlrt.in/gv

๐Ÿš€ *Vulnerability Management Training Demo*
๐Ÿ—“️ *Date:* 28th Oct 2026, 10:00 AM IST
๐Ÿ“*Register Now!:*
https://www.vlrt.in/vr
๐Ÿ‘ฅ *Join Community:*
https://www.vlrt.in/cw
๐Ÿ’ก *Course Content:*
https://www.vlrt.in/vc
▶️ *Demo Videos:*
https://www.vlrt.in/Vm

Monday, 26 February 2024

Q: If Z = 52 and ACT = 48, then what will BAT equal?

 


Q: If Z = 52 and ACT = 48, then what will BAT equal?

  • A) 39

  • B) 46

  • C) 44

  • D) 41

Answer: B) 46

Explanation:

This problem uses a simple letter-to-number substitution code. Here's how it works:

  • Assign Numbers: Each letter of the alphabet is assigned a corresponding number (A=1, B=2, C=3, and so on).

  • Calculate Word Value: Add up the values of each letter in a word to get its numerical equivalent.

  • ACT: 1 + 3 + 20 = 24

  • Since Z has a value of 52 (likely representing 2 x 26), we multiply the word value by 2.

  • ACT value: 24 x 2 = 48

  • Apply to BAT:

  • BAT: 2 + 1 + 20 = 23

  • BAT value: 23 x 2 = 46

Therefore, BAT equals 46.


Hindi


เคช्เคฐเคถ्เคจ: เคฏเคฆि Z = 52 เค”เคฐ ACT = 48, เคคो BAT เค•िเคธเค•े เคฌเคฐाเคฌเคฐ เคนोเค—ा?

  • เค) 39

  • เคฌी) 46

  • เคธी) 44

  • เคกी) 41

เค‰เคค्เคคเคฐ: เคฌी) 46

เคธ्เคชเคท्เคŸीเค•เคฐเคฃ:

เคฏเคน เคธเคฎเคธ्เคฏा เคเค• เคธเคฐเคฒ เค…เค•्เคทเคฐ-เคธे-เคธंเค–्เคฏा เคช्เคฐเคคिเคธ्เคฅाเคชเคจ เค•ोเคก เค•ा เค‰เคชเคฏोเค— เค•เคฐเคคी เคนै। เคฏเคน เคเคธे เค•ाเคฎ เค•เคฐเคคा เคนै:

  • เคธंเค–्เคฏाเคँ เคจिเคฐ्เคฆिเคท्เคŸ เค•เคฐें: เคตเคฐ्เคฃเคฎाเคฒा เค•े เคช्เคฐเคค्เคฏेเค• เค…เค•्เคทเคฐ เค•ो เคเค• เคธंเค—เคค เคธंเค–्เคฏा เคฆी เค—เคˆ เคนै (A=1, B=2, C=3, เค‡เคค्เคฏाเคฆि)।

  • เคถเคฌ्เคฆ เคฎाเคจ เค•ी เค—เคฃเคจा เค•เคฐें: เค•िเคธी เคถเคฌ्เคฆ เค•ा เคธंเค–्เคฏाเคค्เคฎเค• เคธเคฎเค•เค•्เคท เคช्เคฐाเคช्เคค เค•เคฐเคจे เค•े เคฒिเค เค‰เคธเค•े เคช्เคฐเคค्เคฏेเค• เค…เค•्เคทเคฐ เค•ा เคฎाเคจ เคœोเคก़ें।

  • เค…เคงिเคจिเคฏเคฎ: 1 + 3 + 20 = 24

  • เคšूँเค•ि Z เค•ा เคฎाเคจ 52 เคนै (เคธंเคญเคตเคคः 2 x 26 เค•ा เคช्เคฐเคคिเคจिเคงिเคค्เคต เค•เคฐเคคा เคนै), เคนเคฎ เคถเคฌ्เคฆ เคฎाเคจ เค•ो 2 เคธे เค—ुเคฃा เค•เคฐเคคे เคนैं।

  • เค…เคงिเคจिเคฏเคฎ เคฎाเคจ: 24 x 2 = 48

  • เคฌैเคŸ เคชเคฐ เค†เคตेเคฆเคจ เค•เคฐें:

  • เคฌैเคŸ: 2 + 1 + 20 = 23

  • เคฌैเคŸ เคฎाเคจ: 23 x 2 = 46

เค‡เคธเคฒिเค, BAT 46 เค•े เคฌเคฐाเคฌเคฐ เคนै।


Telugu

Q: Z = 52 เฐฎเฐฐిเฐฏు ACT = 48 เฐ…เฐฏిเฐคే, BAT เฐฆేเฐจిเฐ•ి เฐธเฐฎాเฐจం?

  • เฐŽ) 39

  • เฐฌి) 46

  • เฐธి) 44

  • เฐกి) 41

เฐธเฐฎాเฐงాเฐจం: เฐฌి) 46

เฐตిเฐตเฐฐเฐฃ:

เฐˆ เฐธเฐฎเฐธ్เฐฏ เฐธాเฐงాเฐฐเฐฃ เฐ…เฐ•్เฐทเฐฐం เฐจుంเฐกి เฐธంเฐ–్เฐฏ เฐช్เฐฐเฐค్เฐฏాเฐฎ్เฐจాเฐฏ เฐ•ోเฐก్‌เฐจి เฐ‰เฐชเฐฏోเฐ—ిเฐธ్เฐคుంเฐฆి. เฐ‡เฐฆి เฐŽเฐฒా เฐชเฐจి เฐšేเฐธ్เฐคుంเฐฆో เฐ‡เฐ•్เฐ•เฐก เฐ‰ంเฐฆి:

  • เฐธంเฐ–్เฐฏเฐฒเฐจు เฐ•ేเฐŸాเฐฏింเฐšంเฐกి: เฐตเฐฐ్เฐฃเฐฎాเฐฒเฐฒోเฐจి เฐช్เฐฐเฐคి เฐ…เฐ•్เฐทเฐฐాเฐจిเฐ•ి เฐธంเฐฌంเฐงిเฐค เฐธంเฐ–్เฐฏ เฐ•ేเฐŸాเฐฏింเฐšเฐฌเฐกుเฐคుంเฐฆి (A=1, B=2, C=3, เฐฎเฐฐిเฐฏు เฐฎొเฐฆเฐฒైเฐจเฐตి).

  • เฐชเฐฆ เฐตిเฐฒుเฐตเฐจు เฐฒెเฐ•్เฐ•ింเฐšంเฐกి: เฐ’เฐ• เฐชเฐฆంเฐฒోเฐจి เฐช్เฐฐเฐคి เฐ…เฐ•్เฐทเฐฐం เฐฏొเฐ•్เฐ• เฐธంเฐ–్เฐฏా เฐธเฐฎాเฐจเฐคเฐจు เฐชొంเฐฆเฐกాเฐจిเฐ•ి เฐฆాเฐจి เฐตిเฐฒుเฐตเฐฒเฐจు เฐœోเฐกింเฐšంเฐกి.

  • เฐšเฐŸ్เฐŸం: 1 + 3 + 20 = 24

  • Z เฐตిเฐฒుเฐต 52 (เฐฌเฐนుเฐถా 2 x 26เฐจి เฐธూเฐšిเฐธ్เฐคుంเฐฆి) เฐ•เฐฒిเฐ—ి เฐ‰เฐจ్เฐจంเฐฆుเฐจ, เฐฎేเฐฎు เฐชเฐฆ เฐตిเฐฒుเฐตเฐจు 2เฐคో เฐ—ుเฐฃిเฐธ్เฐคాเฐฎు.

  • ACT เฐตిเฐฒుเฐต: 24 x 2 = 48

  • BATเฐ•ి เฐตเฐฐ్เฐคింเฐšంเฐกి:

  • เฐฌ్เฐฏాเฐŸ్: 2 + 1 + 20 = 23

  • BAT เฐตిเฐฒుเฐต: 23 x 2 = 46

เฐ•ాเฐฌเฐŸ్เฐŸి, BAT 46เฐ•ి เฐธเฐฎాเฐจం.


Tamil

 


เฎ•ே: Z = 52 เฎฎเฎฑ்เฎฑுเฎฎ் ACT = 48 เฎŽเฎฉிเฎฒ், BAT เฎŽเฎคเฎฑ்เฎ•ு เฎšเฎฎเฎฎாเฎ• เฎ‡เฎฐுเฎ•்เฎ•ுเฎฎ்?

  • A) 39

  • B) 46

  • C) 44

  • D) 41

เฎชเฎคிเฎฒ்: เฎ†) 46

เฎตிเฎณเฎ•்เฎ•เฎฎ்:

เฎ‡เฎจ்เฎคเฎš் เฎšிเฎ•்เฎ•เฎฒ் เฎ’เฎฐு เฎŽเฎณிเฎฏ เฎŽเฎดுเฎค்เฎคு-เฎŽเฎฃ் เฎฎாเฎฑ்เฎฑுเฎ•் เฎ•ுเฎฑிเฎฏீเฎŸ்เฎŸைเฎช் เฎชเฎฏเฎฉ்เฎชเฎŸுเฎค்เฎคுเฎ•ிเฎฑเฎคு. เฎ‡เฎคு เฎŽเฎต்เฎตாเฎฑு เฎšெเฎฏเฎฒ்เฎชเฎŸுเฎ•ிเฎฑเฎคு เฎŽเฎฉ்เฎชเฎคு เฎ‡เฎ™்เฎ•ே:

  • เฎŽเฎฃ்เฎ•เฎณை เฎ’เฎคுเฎ•்เฎ•ுเฎ™்เฎ•เฎณ்: เฎŽเฎดுเฎค்เฎคுเฎ•்เฎ•เฎณிเฎฉ் เฎ’เฎต்เฎตொเฎฐு เฎŽเฎดுเฎค்เฎคுเฎ•்เฎ•ுเฎฎ் เฎคொเฎŸเฎฐ்เฎชுเฎŸைเฎฏ เฎŽเฎฃ் (A=1, B=2, C=3, เฎฎเฎฑ்เฎฑுเฎฎ் เฎชเฎฒ) เฎ’เฎคுเฎ•்เฎ•เฎช்เฎชเฎŸுเฎฎ்.

  • เฎตாเฎฐ்เฎค்เฎคைเฎฏிเฎฉ் เฎฎเฎคிเฎช்เฎชைเฎ•் เฎ•เฎฃเฎ•்เฎ•ிเฎŸுเฎ™்เฎ•เฎณ்: เฎ’เฎต்เฎตொเฎฐு เฎŽเฎดுเฎค்เฎคிเฎฉ் เฎฎเฎคிเฎช்เฎชுเฎ•เฎณைเฎ•் เฎ•ூเฎŸ்เฎŸி, เฎ…เฎคเฎฉ் เฎŽเฎฃ்เฎฃுเฎ•்เฎ•ுเฎš் เฎšเฎฎเฎฎாเฎฉ เฎฎเฎคிเฎช்เฎชைเฎช் เฎชெเฎฑเฎตுเฎฎ்.

  • เฎšเฎŸ்เฎŸเฎฎ்: 1 + 3 + 20 = 24

  • Z เฎ•்เฎ•ு 52 เฎฎเฎคிเฎช்เฎชு เฎ‡เฎฐுเฎช்เฎชเฎคாเฎฒ் (2 x 26 เฎเฎ•் เฎ•ுเฎฑிเฎ•்เฎ•ுเฎฎ்), เฎจாเฎฎ் เฎตாเฎฐ்เฎค்เฎคை เฎฎเฎคிเฎช்เฎชை 2 เฎ†เฎฒ் เฎชெเฎฐுเฎ•்เฎ•ுเฎ•ிเฎฑோเฎฎ்.

  • ACT เฎฎเฎคிเฎช்เฎชு: 24 x 2 = 48

  • BAT เฎ•்เฎ•ு เฎตிเฎฃ்เฎฃเฎช்เฎชிเฎ•்เฎ•เฎตுเฎฎ்:

  • เฎชேเฎŸ்: 2 + 1 + 20 = 23

  • BAT เฎฎเฎคிเฎช்เฎชு: 23 x 2 = 46

เฎŽเฎฉเฎตே, BAT 46เฎ•்เฎ•ு เฎšเฎฎเฎฎ்.


Spanish

 


P: Si Z = 52 y ACT = 48, ¿a quรฉ serรก BAT?

  • A) 39

  • b) 46

  • c) 44

  • D) 41

Respuesta: B) 46

Explicaciรณn:

Este problema utiliza un cรณdigo simple de sustituciรณn de letra por nรบmero. Asรญ es como funciona:

  • Asignar nรบmeros: a cada letra del alfabeto se le asigna un nรบmero correspondiente (A=1, B=2, C=3, etc.).

  • Calcular el valor de la palabra: sume los valores de cada letra de una palabra para obtener su equivalente numรฉrico.

  • ACTO: 1 + 3 + 20 = 24

  • Dado que Z tiene un valor de 52 (probablemente representa 2 x 26), multiplicamos el valor de la palabra por 2.

  • Valor ACT: 24 x 2 = 48

  • Aplicar a BAT:

  • BATE: 2 + 1 + 20 = 23

  • Valor MTD: 23 x 2 = 46

Por lo tanto, BAT es igual a 46.


French

 


Q : Si Z = 52 et ACT = 48, alors ร  quoi sera รฉgal BAT ?

  • A)39

  • B)46

  • C)44

  • D)41

Rรฉponse : B) 46

Explication:

Ce problรจme utilise un simple code de substitution lettre-chiffre. Voici comment cela fonctionne:

  • Attribuer des numรฉros : chaque lettre de l'alphabet se voit attribuer un numรฉro correspondant (A=1, B=2, C=3, etc.).

  • Calculer la valeur d'un mot : additionnez les valeurs de chaque lettre d'un mot pour obtenir son รฉquivalent numรฉrique.

  • ACTE : 1 + 3 + 20 = 24

  • Puisque Z a une valeur de 52 (reprรฉsentant probablement 2 x 26), nous multiplions la valeur du mot par 2.

  • Valeur ACT : 24 x 2 = 48

  • Appliquer ร  BAT :

  • BATTE : 2 + 1 + 20 = 23

  • Valeur BAT : 23 x 2 = 46

Par consรฉquent, BAT est รฉgal ร  46.



A problem is given to three students whose chances of solving it are 1/2, 1/3, and 1/4 respectively. What is the probability that the problem will be solved?

 Q: A problem is given to three students whose chances of solving it are 1/2, 1/3, and 1/4 respectively. What is the probability that the problem will be solved?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

Answer: C) 3/4

Explanation:

Here's one way to solve this problem:

  1. Probability of NOT solving: Find the probability that none of the students can solve the problem.

  • Student 1: (1 - 1/2) = 1/2

  • Student 2: (1 - 1/3) = 2/3

  • Student 3: (1 - 1/4) = 3/4

  1. Probability of NONE solving: Multiply the individual probabilities of not solving:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probability of AT LEAST ONE solving: Subtract the probability of none solving from 1 (representing certainty):

  • 1 - 1/4 = 3/4

Therefore, the probability that the problem will be solved is 3/4.


Hindi

เคช्เคฐเคถ्เคจ: เคเค• เคธเคฎเคธ्เคฏा เคคीเคจ เค›ाเคค्เคฐों เค•ो เคฆी เค—เคˆ เคนै เคœिเคจเค•े เคนเคฒ เค•เคฐเคจे เค•ी เคธंเคญाเคตเคจा เค•्เคฐเคฎเคถः 1/2, 1/3 เค”เคฐ 1/4 เคนै। เค•्เคฏा เคธंเคญाเคตเคจा เคนै เค•ि เคธเคฎเคธ्เคฏा เคนเคฒ เคนो เคœाเคฏेเค—ी?

  • เค) 7/12

  • เคฌी) 1/2

  • เคธी) 3/4

  • เคกी) 1/4

เค‰เคค्เคคเคฐ: เคธी) 3/4

เคธ्เคชเคท्เคŸीเค•เคฐเคฃ:

เค‡เคธ เคธเคฎเคธ्เคฏा เค•ो เคนเคฒ เค•เคฐเคจे เค•ा เคเค• เคคเคฐीเค•ा เคฏเคนां เคฆिเคฏा เค—เคฏा เคนै:

  1. เคนเคฒ เคจ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เคช्เคฐाเคฏिเค•เคคा เคœ्เคžाเคค เค•ीเคœिเค เค•ि เค•ोเคˆ เคญी เค›ाเคค्เคฐ เคธเคฎเคธ्เคฏा เค•ा เคธเคฎाเคงाเคจ เคจเคนीं เค•เคฐ เคธเค•เคคा।

  • เคตिเคฆ्เคฏाเคฐ्เคฅी 1: (1 - 1/2) = 1/2

  • เค›ाเคค्เคฐ 2: (1 - 1/3) = 2/3

  • เค›ाเคค्เคฐ 3: (1 - 1/4) = 3/4

  1. เค•िเคธी เค•े เคญी เคนเคฒ เคจ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เคนเคฒ เคจ เค•เคฐ เคชाเคจे เค•ी เคต्เคฏเค•्เคคिเค—เคค เคธंเคญाเคตเคจाเค“ं เค•ो เค—ुเคฃा เค•เคฐें:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เค•เคฎ เคธे เค•เคฎ เคเค• เคฌाเคฐ เคนเคฒ เค•เคฐเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा: เค•ोเคˆ เคญी เคนเคฒ เคจ เค•เคฐ เคชाเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा เค•ो 1 เคธे เค˜เคŸाเคँ (เคจिเคถ्เคšिเคคเคคा เค•ा เคช्เคฐเคคिเคจिเคงिเคค्เคต เค•เคฐเคคे เคนुเค):

  • 1 - 1/4 = 3/4

เค…เคคः เคธเคฎเคธ्เคฏा เค•े เคนเคฒ เคนो เคœाเคจे เค•ी เคช्เคฐाเคฏिเค•เคคा 3/4 เคนै।


Telugu

 เฐช్เฐฐ: เฐฎుเฐ—్เฐ—ుเฐฐు เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅుเฐฒเฐ•ు เฐ’เฐ• เฐธเฐฎเฐธ్เฐฏ เฐ‡เฐต్เฐตเฐฌเฐกింเฐฆి, เฐฆాเฐจి เฐชเฐฐిเฐท్เฐ•ాเฐฐ เฐ…เฐตเฐ•ాเฐถాเฐฒు เฐตเฐฐుเฐธเฐ—ా 1/2, 1/3 เฐฎเฐฐిเฐฏు 1/4. เฐธเฐฎเฐธ్เฐฏ เฐชเฐฐిเฐท్เฐ•ాเฐฐเฐฎเฐฏ్เฐฏే เฐธంเฐญాเฐต్เฐฏเฐค เฐเฐฎిเฐŸి?

  • เฐŽ) 7/12

  • เฐฌి) 1/2

  • เฐธి) 3/4

  • เฐกి) 1/4

เฐธเฐฎాเฐงాเฐจం: เฐธి) 3/4

เฐตిเฐตเฐฐเฐฃ:

เฐˆ เฐธเฐฎเฐธ్เฐฏเฐจు เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐกాเฐจిเฐ•ి เฐ‡เฐ•్เฐ•เฐก เฐ’เฐ• เฐฎాเฐฐ్เฐ—ం เฐ‰ంเฐฆి:

  1. เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐค: เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅుเฐฒెเฐตเฐฐూ เฐธเฐฎเฐธ్เฐฏเฐจు เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐฒేเฐจి เฐธంเฐญాเฐต్เฐฏเฐคเฐจు เฐ•เฐจుเฐ—ొเฐจంเฐกి.

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 1: (1 - 1/2) = 1/2

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 2: (1 - 1/3) = 2/3

  • เฐตిเฐฆ్เฐฏాเฐฐ్เฐฅి 3: (1 - 1/4) = 3/4

  1. เฐเฐฆీ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐค: เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐต్เฐฏเฐ•్เฐคిเฐ—เฐค เฐธంเฐญాเฐต్เฐฏเฐคเฐฒเฐจు เฐ—ుเฐฃింเฐšంเฐกి:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เฐ•เฐจీเฐธం เฐ’เฐ• เฐชเฐฐిเฐท్เฐ•ాเฐฐం เฐฏొเฐ•్เฐ• เฐธంเฐญాเฐต్เฐฏเฐค: เฐเฐฆీ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐจి เฐธంเฐญాเฐต్เฐฏเฐคเฐจు 1 เฐจుంเฐกి เฐคీเฐธిเฐตేเฐฏంเฐกి (เฐจిเฐถ్เฐšเฐฏเฐคเฐจు เฐธూเฐšిเฐธ్เฐคుంเฐฆి):

  • 1 - 1/4 = 3/4

เฐ…ంเฐฆుเฐตเฐฒ్เฐฒ, เฐธเฐฎเฐธ్เฐฏ เฐชเฐฐిเฐท్เฐ•เฐฐింเฐšเฐฌเฐกే เฐธంเฐญాเฐต్เฐฏเฐค 3/4.


Tamil

 เฎ•ே: เฎ’เฎฐு เฎšிเฎ•்เฎ•เฎฒைเฎค் เฎคீเฎฐ்เฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎตாเฎฏ்เฎช்เฎชு เฎฎுเฎฑைเฎฏே 1/2, 1/3 เฎฎเฎฑ்เฎฑுเฎฎ் 1/4 เฎ†เฎ•ிเฎฏ เฎฎூเฎฉ்เฎฑு เฎฎாเฎฃเฎตเฎฐ்เฎ•เฎณுเฎ•்เฎ•ு เฎตเฎดเฎ™்เฎ•เฎช்เฎชเฎŸுเฎ•ிเฎฑเฎคு. เฎชிเฎฐเฎš்เฎšเฎฉை เฎคீเฎฐுเฎฎ் เฎจிเฎ•เฎด்เฎคเฎ•เฎตு เฎŽเฎฉ்เฎฉ?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

เฎชเฎคிเฎฒ்: C) 3/4

เฎตிเฎณเฎ•்เฎ•เฎฎ்:

เฎ‡เฎจ்เฎค เฎšிเฎ•்เฎ•เฎฒை เฎคீเฎฐ்เฎ•்เฎ• เฎ’เฎฐு เฎตเฎดி เฎ‡เฎ™்เฎ•ே:

  1. เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎฎเฎฒ் เฎ‡เฎฐுเฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎฎாเฎฃเฎตเฎฐ்เฎ•เฎณ் เฎฏாเฎฐுเฎฎ் เฎšிเฎ•்เฎ•เฎฒைเฎค் เฎคீเฎฐ்เฎ•்เฎ• เฎฎுเฎŸிเฎฏாเฎค เฎจிเฎ•เฎด்เฎคเฎ•เฎตைเฎ•் เฎ•เฎฃ்เฎŸเฎฑிเฎฏเฎตுเฎฎ்.

  • เฎฎாเฎฃเฎตเฎฐ் 1: (1 - 1/2) = 1/2

  • เฎฎாเฎฃเฎตเฎฐ் 2: (1 - 1/3) = 2/3

  • เฎฎாเฎฃเฎตเฎฐ் 3: (1 - 1/4) = 3/4

  1. เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎค เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎคீเฎฐ்เฎ•்เฎ•ாเฎค เฎคเฎฉிเฎช்เฎชเฎŸ்เฎŸ เฎจிเฎ•เฎด்เฎคเฎ•เฎตுเฎ•เฎณை เฎชெเฎฐுเฎ•்เฎ•เฎตுเฎฎ்:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. เฎ•ுเฎฑைเฎจ்เฎคเฎชเฎŸ்เฎšเฎฎ் เฎ’เฎฐு เฎคீเฎฐ்เฎตுเฎ•்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு: เฎŽเฎคுเฎตுเฎฎ் เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸாเฎฎเฎฒ் เฎ‡เฎฐுเฎช்เฎชเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตை 1 เฎ‡เฎฒிเฎฐுเฎจ்เฎคு เฎ•เฎดிเฎ•்เฎ•เฎตுเฎฎ் (เฎ‰เฎฑுเฎคிเฎฏாเฎฉ เฎคเฎฉ்เฎฎைเฎฏைเฎ•் เฎ•ுเฎฑிเฎ•்เฎ•ுเฎฎ்):

  • 1 - 1/4 = 3/4

เฎŽเฎฉเฎตே, เฎšிเฎ•்เฎ•เฎฒ் เฎคீเฎฐ்เฎ•்เฎ•เฎช்เฎชเฎŸுเฎตเฎคเฎฑ்เฎ•ாเฎฉ เฎจிเฎ•เฎด்เฎคเฎ•เฎตு 3/4 เฎ†เฎ•ுเฎฎ்.


Spanish

 P: Se les da un problema a tres estudiantes cuyas posibilidades de resolverlo son 1/2, 1/3 y 1/4 respectivamente. ¿Cuรกl es la probabilidad de que se resuelva el problema?

  • A) 7/12

  • b) 1/2

  • c) 3/4

  • D) 1/4

Respuesta: C) 3/4

Explicaciรณn:

Aquรญ hay una forma de resolver este problema:

  1. Probabilidad de NO resolver: Encuentre la probabilidad de que ninguno de los estudiantes pueda resolver el problema.

  • Estudiante 1: (1 - 1/2) = 1/2

  • Estudiante 2: (1 - 1/3) = 2/3

  • Estudiante 3: (1 - 1/4) = 3/4

  1. Probabilidad de que NINGUNO resuelva: Multiplique las probabilidades individuales de no resolver:

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probabilidad de que AL MENOS UNO resuelva: Reste la probabilidad de que ninguno resuelva de 1 (que representa certeza):

  • 1 - 1/4 = 3/4

Por tanto, la probabilidad de que el problema se resuelva es 3/4.


French

 Q : Un problรจme est confiรฉ ร  trois รฉlรจves dont les chances de le rรฉsoudre sont respectivement de 1/2, 1/3 et 1/4. Quelle est la probabilitรฉ que le problรจme soit rรฉsolu ?

  • A) 7/12

  • B) 1/2

  • C) 3/4

  • D) 1/4

Rรฉponse : C) 3/4

Explication:

Voici une faรงon de rรฉsoudre ce problรจme :

  1. Probabilitรฉ de NE PAS rรฉsoudre : Trouvez la probabilitรฉ qu’aucun des รฉlรจves ne puisse rรฉsoudre le problรจme.

  • ร‰tudiant 1 : (1 - 1/2) = 1/2

  • ร‰tudiant 2 : (1 - 1/3) = 2/3

  • ร‰lรจve 3 : (1 - 1/4) = 3/4

  1. Probabilitรฉ de ne pas rรฉsoudre : Multipliez les probabilitรฉs individuelles de ne pas rรฉsoudre :

  • (1/2) * (2/3) * (3/4) = 1/4

  1. Probabilitรฉ d’AU MOINS UNE rรฉsolution : soustrayez la probabilitรฉ qu’aucune solution ne soit rรฉsolue de 1 (reprรฉsentant la certitude) :

  • 1 - 1/4 = 3/4

La probabilitรฉ que le problรจme soit rรฉsolu est donc de 3/4.